# Flatbands in tight-binding lattices with anisotropic potentials

### v1

@@ Appendix B Duality transformation for the quasiperiodic potential
$E(k)\psi_{z}(k)=\lambda\sin(\beta z)\psi_{z}(k)$ (31)
$-2\cos\left(\frac{k}{2}\right)\left[e^{\frac{ik}{2}}\psi_{z+1}(k)+e^{-\frac{ik}{2}}\psi_{z-1}(k)\right]\;.$
with |$\ket{\psi(k)}=\sum_{z}\psi_{z}(k)\ket{z}$. This eigenproblem is equivalent to that of the Hamiltonian in Eq. (55) of the Ref. 32 up to the additional phase factors $e^{\pm\frac{ik}{2}}$ in the hopping parameters. Inspired by the duality transformations used in Refs. 42, 32 we use a modified Fourier transformation connecting $z$-space to $m$-space
$\psi_{z}(k)=\frac{1}{\sqrt{L}}e^{i\gamma z}\sum_{m}g_{m}e^{im\beta z}(i)^{m}e^{\frac{imk}{2}}\;.$ (32)
The value of parameter $\gamma$ is independent of $m$ and $\beta$, and is to be determined later. $L$ is the size of the lattice along the $z$ direction. Plugging the modified Fourier transform into Eq. (31) we arrive at
@@
$-4\cos\left(\frac{k}{2}\right)\cos\left(m\beta+\gamma+k/2\right)g_{m}\;.$ (33)
Setting $\gamma+k/2=\pi/2$ so that $\cos\left(m\beta+\gamma+k/2\right)$ $=-\sin(m\beta)$, we arrive at the dual problem:
$E(k)g_{m}=-\frac{\lambda}{2}\left[g_{m-1}e^{-\frac{ik}{2}}+g_{m+1}e^{\frac{ik}{2}}\right]+$
$+4\cos\left(\frac{k}{2}\right)\sin(m\beta)g_{m}\;.$ (34)
This eigenproblem is equivalent to Eq. (31) up to a swap of $\lambda$ and $4\cos\left(\frac{k}{2}\right)$. Therefore the Hamiltonian (12) with potential (17) has a duality transformation similar to that of the Aubry-André model, but for every individual momentum $k$. We expect a localization to delocalization transition happen at a self-dual point
@@
This expression gives a localization-delocalization transition curve in the $(\lambda,k)$-space and we can determine the boundary in $k$:
@@
For a fixed $\lambda\leq 4$, all eigenstates, including the $E=0$ flatband eigenstates, with energies $E\in\{E(k):\absolutevalue{k}<k_{c}\}$ delocalize along the $z$ direction, while all eigenstates with $E\in\{E(k):\absolutevalue{k}>k_{c}\}$ are localized along the $z$ direction. The Fig. 13 depicts the mobility edges in the $k$-space at fixed $\lambda$.
Figure 13: Eigenvalues as a function of momentum for Hamiltonian (12) for potential (17) with (a) $\lambda=2$ and (b) $\lambda=3$. Red (blue) color marked the region with localized (delocalized) eigenfunctions along the $z$ direction. The color marked is based on Eq. (36). Note that increasing $\lambda$ shrinks the delocalized regimes.

### v2

@@ Appendix B Duality transformation for the quasiperiodic potential
$E(k)\psi_{z}(k)=\lambda\sin(\beta z)\psi_{z}(k)$
$-2\cos\left(\frac{k}{2}\right)\left[e^{\frac{ik}{2}}\psi_{z+1}(k)+e^{-\frac{ik}{2}}\psi_{z-1}(k)\right],$ (31)
with |$\ket{\psi(k)}=\sum_{z}\psi_{z}(k)\ket{z}$. This eigenproblem is equivalent to that of the Hamiltonian in Eq. (55) of the Ref. [32] up to the additional phase factors $e^{\pm\frac{ik}{2}}$ in the hopping parameters. Inspired by the duality transformations used in Refs. [45, 32], we use a modified Fourier transformation connecting $z$ space to $m$ space:
$\psi_{z}(k)=\frac{1}{\sqrt{L}}e^{i\gamma z}\sum_{m}g_{m}e^{2\pi im\beta z}(i)^{m}e^{\frac{imk}{2}}\;.$ (32)
The value of parameter $\gamma$ is independent of $m$ and $\beta$ and is to be determined later. $L$ is the size of the lattice along the $z$ direction. Plugging the modified Fourier transform into Eq. (B) we arrive at
@@
$-4\cos\left(\frac{k}{2}\right)\cos\left(2\pi m\beta+\gamma+k/2\right)g_{m}\;.$ (33)
Setting $\gamma+k/2=\pi/2$, so $\cos\left(2\pi m\beta+\gamma+k/2\right)$ $=-\sin(2\pi m\beta)$, we arrive at the dual problem:
$+4\cos\left(\frac{k}{2}\right)\sin(2\pi m\beta)g_{m}\;.$ (34)
This eigenproblem is equivalent to Eq. (B) up to a swap of $\lambda$ and $4\cos\left(\frac{k}{2}\right)$. Therefore, the Hamiltonian (12) with potential (17) has a duality transformation similar to that of the Aubry-André model, but for every individual momentum $k$. We expect a localization to delocalization transition happen at a self-dual point:
@@
This expression gives a localization-delocalization transition curve in the $(\lambda,k)$ space and we can determine the boundary in $k$:
@@
For a fixed $\lambda\leq 4$, all eigenstates, including the $E=0$ flat-band eigenstates, with energies $E\in\{E(k):\absolutevalue{k}<k_{c}\}$ delocalize along the $z$ direction, while all eigenstates with $E\in\{E(k):\absolutevalue{k}>k_{c}\}$ are localized along the $z$ direction. Figure 13 depicts the mobility edges in the $k$ space at fixed $\lambda$.
Figure 13: Eigenvalues as a function of momentum for Hamiltonian (12) for potential (17) with (a) $\lambda=2$ and (b) $\lambda=3$. Red (blue) marks the region with localized (delocalized) eigenfunctions along the $z$ direction. The color marked is based on Eq. (36). Note that increasing $\lambda$ shrinks the delocalized regimes.